# Diffusion differentiable resampling

### v1

@@ Appendix F Gaussian mixture resampling
where $x\in\mathbb{R}^{d}$, $y\in\mathbb{R}$, $d=8$, and the number of components $c=5$. The posterior distribution $\pi(x)\propto\phi(x)\,p(y{\;|\;}x)$ is also a Gaussian mixture $\pi(x)=\sum_{i=1}^{c}\Omega_{i}\,\mathrm{N}(x;\mathcal{M}_{i},\mathcal{V}_{i})$ (Zhao2025B0SMC), given by
$\begin{split}G_{i}&=H\,v_{i}\,H^{\mkern-1.5mu\mathsf{T}}+\Xi,\\ \overline{\Omega}_{i}&=\omega_{i}\,\mathrm{N}(y;H\,m_{i},G_{i}),\\ \Omega_{i}&=\overline{\Omega}_{i}\,/\,\sum_{j=1}^{c}\overline{\Omega}_{j},\\ \mathcal{M}_{i}&=m_{i}+v_{i}\,H^{\mkern-1.5mu\mathsf{T}}\,G^{-1}_{i}\,(y-H\,m_{i}),\\ \mathcal{V}_{i}&=v_{i}-v_{i}\,H^{\mkern-1.5mu\mathsf{T}}\,G^{-1}_{i}\,H^{\mkern-1.5mu\mathsf{T}}\,v_{i}.\end{split}$
@@
In literature, the resampling variance is usually defined as the variance of the resampling algorithm itself conditioned on the input samples, that is, $\operatorname{Var}\Bigl[\frac{1}{N}\sum_{i=1}^{N}X_{i}^{*}{\;|\;}\{(w_{i},X_{i})\}_{i=1}^{N}\Bigr]$. This is useful to heuristically gauge the noise level of resampling. However, this does not inform how well the re-samples approximate the true distribution $\pi$. With a slight abuse of terminology, we define the resampling variance as the mean estimator error $\operatorname{\mathbb{E}}\Bigl[\bigl(\frac{1}{N}\sum_{i=1}^{N}X_{i}^{*}-\mathcal{M}\bigr)^{2}\Bigr]$, where $\mathcal{M}$ stands for the true mean of the Gaussian mixture posterior $\pi$. The expectation is approximateed by the 100 independent Monte Carlo runs.
@@ Appendix H Linear Gaussian SSM
The optimiser L-BFGS is implemented using JAXopt (Blondel2021) scipyminimize wrapper with default settings and initial parameters $\theta+1$.
We observe a numerical issue due to the gradient estimate by the resampling methods. The L-BFGS optimiser frequently diverges. This is expected, since L-BFGS is highly sensitive to the quality of gradient estimate (Xie2020); see also Zhao2023MFS for an empirical validation. As such, when we report the results we have defined a convergent run by 1) positive convergence flag returned by the L-BFGS optimiser and 2) the error $\lVert\theta-\widehat{\theta}\rVert_{2}<1.9$ is not absurd. It turns out that the effect is especially pronounced for the soft and Gumbel-Softmax resampling, where at 100 runs, only nearly 10% return a successful flag of the optimiser, and all their estimated parameters diverge to a meaningless position. This was not problematic for the diffusion and OT samplers, where they have approximately 80% success rate in average.
@@
Let us focus on the diffusion resampling in Tables 7 and 9. In Tables 7 we find that the loss function estimation error seems to be invariant in the time $T$ and steps $K$ when using an ODE solver (except for Lord–Rougemont), giving a relatively large error $2.72$. With a fixed $T$, it is not clear if increasing $K$ will improve the log-likelihood estimate. This also resonates with Table 9 that increasing $K$ can possibly lead to worst even exploding parameter estimation. This may be caused by numerical errors when back-propagating gradients through SDE solvers which can be addressed by, for instance, Kidger2021gradient. On the other hand, with $K$ fixed, increasing $T$ improves the result. Like in the previous experiments, the Jentzen–Kloeden exponential integrator consistently achieves the best result, albeit marginally, while Lord–Rougemont is the most sensitive to $K$.
@@
Gumbel (0.1) $2.79\pm 2.14$ $4.83\pm 5.76$ $14.14\pm 0.00$
Gumbel (0.3) $2.75\pm 2.17$ $4.89\pm 5.42$ $14.14\pm 0.00$
Gumbel (0.5) $2.73\pm 2.22$ $5.10\pm 5.93$ $14.14\pm 0.00$
Soft (0.5) $3.63\pm 2.10$ $7.75\pm 11.19$ $14.14\pm 0.00$
Soft (0.7) $3.08\pm 1.88$ $5.34\pm 7.82$ $14.14\pm 0.00$
Soft (0.9) $2.85\pm 1.80$ $4.66\pm 5.68$ $14.14\pm 0.01$
Multinomial $2.80\pm 1.84$ $5.49\pm 6.87$ $14.14\pm 0.00$
@@ Appendix I Prey-predator model
We simulate the model with Milstein’s method and generate data at $t\in[0,3]$ with 256 dicretisation steps. The discretisation results in an SSM $Z_{j+1}=f(Z_{j},\epsilon_{j})$, where $Z_{j}\in\mathbb{R}^{2}$ encodes $C(t_{j})$ and $R(t_{j})$, and $\epsilon_{j}\in\mathbb{R}^{2}$ stands for Brownian motion increment. We train a neural network to approximate the SDE dynamics by minimising the negative log-likelihood, estimated by a particle filter with resampling. The neural network is simply a two-layers fully connected network with residual connection, see Figure 7, mimicking a discrete SDE solver. No prior knowledge of the dynamics is incorporated into the neural network. We train the neural network at a fixed 1,000 iterations by a Lion optimiser (Chen2023) with learning rate 0.005. At testing, we make 100 independent predictions from the learnt model and compute the root mean square error (RMSE) with respect to a reference trajectory generated using the true dynamics. We use 64 particles, and repeat the experiments for 20 times.
@@
We assume that the pendulum dynamics evolve according to Equation (16), with parameters $g=9.81$ and $l=0.4$. Using the known dynamics and the observation model specified in Equation (16), we generate a single sequence of observational data over $t\in[0,4]$ using 256 discretisation steps (which is $j=0,1,\ldots,256$ in discrete time) and starting from the initial state $Z_{0}=\begin{bmatrix}\pi\,/\,4&0\end{bmatrix}$. The observation noise $\sigma_{\mathrm{obs}}=0.01$ and process noise intensity $\sigma_{\zeta}^{2}=0.01$ are fixed and assumed to be known. We jointly optimise the two neural networks, $f_{\theta}$ and $r_{\phi}$, parameterising the transition function and the decoder, respectively, by minimising the negative-log likelihood estimated by the particle filter. As detailed in the main text, we use resampling at each filtering step. For numerical and optimisation stability, we use $\sigma_{\mathrm{train}}=10\,\sigma_{\mathrm{obs}}$ and scale the likelihood formulation. In particular, we compute the log-likelihood as the mean over pixel dimensions (rather than the sum) and rescale the result by a factor $\lambda^{-1}$ (where $\lambda=10^{4}$). We train the neural networks for 3000 iterations. The parameters are updated using the Adam optimiser (learning rate $10^{-4}$) with global gradient clipping (maximum norm 1.0).
@@
We evaluate the learnt models by comparing the average SSIM and PSNR of image sequences obtained by unrolling the dynamics from $Z_{0}$ using the learnt transition function $f_{\theta}$, and then generating images by passing the resulting state trajectory through the learnt decoder $r_{\phi}$. All results are averaged over $5$ independent runs, using $N=32$ particles to approximate the negative log-likelihood during training. Results for different configurations of the diffusion resampler are presented in Table 13. Results for different configurations of the baselines are presented in Table 14. Figure 8 (left) shows the median loss evolution during training for the best model configuration (in terms of mean SSIM and PSNR) in each resampling class. As can be seen, diffusion resampling achieves the lowest loss.
@@
Because both the transition model $f_{\theta}$ and the decoder $r_{\phi}$ are learnt only through their ability to reproduce the observed image sequence, the latent state is in general identifiable only up to (possibly nonlinear) invertible transformations, and there is no unique correspondence between the learnt states and the true physical coordinates. For this reason, we do not directly evaluate the pendulum dynamics in state space, but instead assess model quality in the observation space via image reconstruction metrics. A similar evaluation strategy is common in recent work on latent SDEs for high-dimensional sequence data. For example, both Bartosh2025sde and Course2023amortized primarily report qualitative comparisons of generated image sequences in comparable settings, rather than quantitative metrics in the latent state space, reflecting the view that high-quality reconstruction is indicative of a meaningful latent representation. Nevertheless, for the diffusion resampler we find that, in some runs, the learnt dynamics closely match the true pendulum dynamics up to a linear transformation of the latent coordinates. Figure 8 (right) illustrates one such example with the latent dynamics learnt using diffusion resampling, where a simple linear transformation of the latent trajectory yields a close match to the ground-truth pendulum state (coefficient of determination $R^{2}\approx 0.982$). This indicates that the model has learnt a meaningful representation of the underlying pendulum dynamics, albeit in a transformed latent coordinate system.
@@
Table 13: Prediction quality in terms of structural similarity index measure (SSIM) and PSNR (higher the better) for the pendulum dynamics tracking experiment, with respect to models trained using diffusion resampling. Each result is presented in terms of mean over 5 independent runs, together with the corresponding standard deviation, recorded after 3000 training iterations. For two configurations only 1 out of 5 (*) and 3 out of 5 (**) runs, respectively, were non-divergent.
@@
Table 14: Prediction quality (SSIM, PSNR) for the pendulum dynamics tracking experiment, with respect to models trained using OT, Gumbel and Soft resampling.
@@
Both the neural network modelling the latent dynamics, $f_{\theta}$, and the neural network modelling the decoder, $r_{\phi}$, operate on a feature vector $\mathbf{h}(Z_{j})=\begin{bmatrix}\sin(Z_{j}^{(1)})&\cos(Z_{j}^{(1)})&Z_{j}^{(2)}\,/\,10\end{bmatrix}$, which embeds the angle and scales the velocity. While the system is initialised with physical coordinates $Z_{0}$, both networks are unconstrained and unregularised for all $t>0$. As discussed above, this means that the system is free to learn a latent manifold sufficient for image reconstruction, even if the latent representation is a non-linear transformation of the true physical coordinate system. We employ a SIREN (Sitzmann2019siren) architecture to model the latent dynamics. The network takes the concatenated input $\begin{bmatrix}\mathbf{h}(Z_{j})&\zeta_{j}\end{bmatrix}\in\mathbb{R}^{5}$ and outputs the state derivative. The update rule is $Z_{j+1}=Z_{j}+\mathrm{Net}(\begin{bmatrix}\mathbf{h}(Z_{j})&\zeta_{j}\end{bmatrix})\,\Delta_{j}$. The network consists of 3 hidden layers of 256 units with sine activations ($\omega_{0}=8.0$ in the first layer). The decoder maps $\mathbf{h}(Z_{j})$ to observations. It first processes the input via two linear layers (16 and 256 units), reshaping the output into a spatial feature map of size $4\times 4\times 16$. This is followed by three transposed convolution layers (kernel size $3\times 3$, stride 2) with output channels of 32, 16, and 1, respectively, to upsample to the final $32\times 32$ image. ReLU activations are used for all hidden layers, while the output layer is linear.
@@
Figure 8: (Left) Median loss during the training process for the top performing configuration in each resampling class. (Right) Ground-truth pendulum state trajectories (dashed) and predictions in the original neural (solid with crosses) and transformed (solid) coordinate systems, for a top-performing model trained with diffusion resampling ($T=1$, $K=4$). The dynamics model is trained until convergence, and correspond to the image predictions presented in Figure 6.
Appendix K Take-away messages
@@
  • Diffusion resampling is useful not only for differentiation, but also for reducing resampling error. The primary reason is that diffusion resampling can take more information (e.g., $\pi_{\mathrm{ref}}$) of the target into account. When we have a complete information $\pi_{\mathrm{ref}}=\pi$, the diffusion resampling is an optimal resampling algorithm. Corollary 1 explicitly shows how the resampling error can be controlled depending on how well $\pi_{\mathrm{ref}}$ approximates $\pi$, calibrated by $b$. In contrast, multinomial resampling only uses information from the given samples $\{(w_{i},X_{i})\}_{i=1}^{N}$ which uses incomplete and less information compared to diffusion resampling based on first principle.
@@
  • If a good reference distribution is hard to construct, diffusion resampling may not be optimal for a pure resampling problem outside the SSM context. For instance, as shown in the Gaussian mixture experiment, the large discrepancy between the Gaussian reference and the highly non-Gaussian target may prolong the diffusion time. However, for SSMs, thanks to their sequential structure, constructing an effective reference is more straightforward.
Table 15: Comparison among commonly used resampling schemes with their calibration parameters. The computational complexity is analysed per sample which can be embarrassingly parallelised over the samples. The complexity of OT is given by Luo2023OT, where we here parametrise the Sinkhorn precision with the regularisation parameter $\varepsilon$ to unify comparison. By “fully differentiable” we mean that the pathwise gradient is well defined, for instance, the soft resampling is only partially differentiable.
@@
Diffusion $T$ Yes Yes, as $N(T)\to\infty$ No O(TN)O(T\,N)
OT $\varepsilon$ Yes Yes, as $N(\varepsilon)\to\infty$ No $O(N^{2}\log(N^{-1})\,/\,\varepsilon)$

### v2

@@ Appendix F Gaussian mixture resampling
where $x\in\mathbb{R}^{d}$, $y\in\mathbb{R}$, $d=8$, and the number of components $c=5$. The posterior distribution $\pi(x)\propto\phi(x)\,p(y{\;|\;}x)$ is also a Gaussian mixture $\pi(x)=\sum_{i=1}^{c}\Omega_{i}\,\mathrm{N}(x;\mathcal{M}_{i},\mathcal{V}_{i})$ (Zhao, 2025), given by
$\begin{split}G_{i}&=H\,v_{i}\,H^{\mkern-1.5mu\mathsf{T}}+\Xi,\\ \overline{\Omega}_{i}&=\omega_{i}\,\mathrm{N}(y;H\,m_{i},G_{i}),\\ \Omega_{i}&=\overline{\Omega}_{i}\,/\,\sum_{j=1}^{c}\overline{\Omega}_{j},\\ \mathcal{M}_{i}&=m_{i}+v_{i}\,H^{\mkern-1.5mu\mathsf{T}}\,G^{-1}_{i}\,(y-H\,m_{i}),\\ \mathcal{V}_{i}&=v_{i}-v_{i}\,H^{\mkern-1.5mu\mathsf{T}}\,G^{-1}_{i}\,H\,v_{i}.\end{split}$
@@
In literature, the resampling variance is usually defined as the variance of the resampling algorithm itself conditioned on the input samples, that is, $\operatorname{Var}\Bigl[\frac{1}{N}\sum_{i=1}^{N}X_{i}^{*}{\;|\;}\{(w_{i},X_{i})\}_{i=1}^{N}\Bigr]$. This is useful to heuristically gauge the noise level of resampling. However, this does not inform how well the re-samples approximate the true distribution $\pi$. With a slight abuse of terminology, we define the resampling variance as the mean estimator error $\operatorname{\mathbb{E}}\Bigl[\bigl(\frac{1}{N}\sum_{i=1}^{N}X_{i}^{*}-\mathcal{M}\bigr)^{2}\Bigr]$, where $\mathcal{M}$ stands for the true mean of the Gaussian mixture posterior $\pi$. The expectation is approximated by the 100 independent Monte Carlo runs.
@@ Appendix H Linear Gaussian SSM
The optimiser L-BFGS is implemented using JAXopt (Blondel et al., 2021) scipyminimize wrapper with default settings and initial parameters $\theta+1$.
We observe a numerical issue due to the gradient estimate by the resampling methods. The L-BFGS optimiser frequently diverges. This is expected, since L-BFGS is highly sensitive to the quality of gradient estimate (Xie et al., 2020); see also Zhao and Sarmarvuori (2023) for an empirical validation. As such, when we report the results we have defined a convergent run by 1) positive convergence flag returned by the L-BFGS optimiser and 2) the error $\lVert\theta-\widehat{\theta}\rVert_{2}<1.9$ is not absurd. It turns out that the effect is especially pronounced for the soft and Gumbel-Softmax resampling, where at 100 runs, only nearly 10% return a successful flag of the optimiser, and all their estimated parameters diverge to a meaningless position. This was not problematic for the diffusion and OT samplers, where they have approximately 80% success rate in average.
@@
Let us focus on the diffusion resampling in Tables 7 and 9. In Tables 7 we find that the loss function estimation error seems to be invariant in the time $T$ and steps $K$ when using an ODE solver (except for Lord–Rougemont), giving a relatively large error $2.72$. With a fixed $T$, it is not clear if increasing $K$ will improve the log-likelihood estimate. This also resonates with Table 9 that increasing $K$ can possibly lead to worst even exploding parameter estimation. This may be caused by numerical errors when back-propagating gradients through SDE solvers which can be addressed by, for instance, Kidger et al. (2021). On the other hand, with $K$ fixed, increasing $T$ improves the result. Like in the previous experiments, the Jentzen–Kloeden exponential integrator consistently achieves the best result, albeit marginally, while Lord–Rougemont is the most sensitive to $K$.
@@
Gumbel (0.1) $2.79\pm 2.14$ $4.83\pm 5.76$ NaN
Gumbel (0.3) $2.75\pm 2.17$ $4.89\pm 5.42$ NaN
Gumbel (0.5) $2.73\pm 2.22$ $5.10\pm 5.93$ NaN
Soft (0.5) $3.63\pm 2.10$ $7.75\pm 11.19$ NaN
Soft (0.7) $3.08\pm 1.88$ $5.34\pm 7.82$ NaN
Soft (0.9) $2.85\pm 1.80$ $4.66\pm 5.68$ NaN
Multinomial $2.80\pm 1.84$ $5.49\pm 6.87$ NaN
@@ Appendix I Prey-predator model
We simulate the model with Milstein’s method and generate data at $t\in[0,3]$ with 256 dicretisation steps. The discretisation results in an SSM $Z_{j+1}=f(Z_{j},\epsilon_{j})$, where $Z_{j}\in\mathbb{R}^{2}$ encodes $C(t_{j})$ and $R(t_{j})$, and $\epsilon_{j}\in\mathbb{R}^{2}$ stands for Brownian motion increment. We train a neural network to approximate the SDE dynamics by minimising the negative log-likelihood, estimated by a particle filter with resampling. The neural network is simply a two-layers fully connected network with residual connection, see Figure 7, mimicking a discrete SDE solver. No prior knowledge of the dynamics is incorporated into the neural network. We train the neural network at a fixed 1,000 iterations by a Lion optimiser (Chen et al., 2023b) with learning rate 0.005. At testing, we make 100 independent predictions from the learnt model and compute the root mean square error (RMSE) with respect to a reference trajectory generated using the true dynamics. We use 64 particles, and repeat the experiments 20 times.
@@
We assume that the pendulum dynamics evolve according to Equation (16), with parameters $g=9.81$ and $l=0.4$. Using the known dynamics and the observation model specified in Equation (16) with observation noise $\sigma_{\mathrm{obs}}=0.01$, we generate a single sequence of observational data over $t\in[0,4]$ using 256 discretisation steps (which is $j=0,1,\ldots,256$ in discrete time) and starting from the initial state $Z_{0}=\begin{bmatrix}\pi\,/\,4&0\end{bmatrix}$. During training, we jointly optimise the two neural networks, $f_{\theta}$ and $r_{\phi}$, parameterising the transition function and the decoder, respectively, by minimising the negative log-likelihood (NLL) estimated by the particle filter.
For completeness, we consider two experiment settings: in the first setting we employ effective sample size (ESS)-based adaptive resampling under a tempered observation likelihood. Although the tempering modifies the original model, it is a common practice in deep learning for training complex models (see, e.g., Ng et al., 2025). Here, the process noise covariance is fixed by $\Lambda_{\zeta}=\sigma_{\zeta}^{2}\,I_{2}$, where $\sigma_{\zeta}^{2}=0.01$. We use $\sigma_{\mathrm{train}}=10\,\sigma_{\mathrm{obs}}$ and scale the observation log-likelihood. In particular, we compute the log-likelihood as the mean over pixel dimensions (rather than the sum) and rescale the result by a factor $\kappa^{-1}$ (where $\kappa=10^{4}$). We use a resampling threshold such that we resample if the ESS is less than $N=32$. We train the neural networks for 3000 iterations. The learnable parameters are updated using the Adam optimiser (learning rate $10^{-4}$) with global gradient clipping (maximum norm 1.0) and $\beta_{2}=0.999$ (other Adam defaults as in Optax). We use $N=32$ particles to approximate the negative log-likelihood during training.
In the second setting, we consider a more challenging regime in which we force resampling at every filtering step after $j>10$. Here, we use process noise with $\Lambda_{\zeta}=\sigma_{\zeta}^{2}\,\operatorname{diag}(0,1)$ with $\sigma_{\zeta}^{2}=0.1$, and make an additive noise assumption in the learned transition model. We use $\sigma_{\mathrm{train}}=5\,\sigma_{\mathrm{obs}}$ and no other scaling of the log-likelihood compared to the true observation model. We train the neural networks for 1500 iterations. The learnable parameters are updated using the Adam optimiser (learning rate $2\times 10^{-4}$) with global gradient clipping (maximum norm 1.0) and $\beta_{2}=0.99$ (other Adam defaults as in Optax). Here, we use $N=16$ particles. The optimiser omits at most 10 consecutive non-finite parameter updates by leaving the current parameter values unchanged.
Results and evaluation
@@
We evaluate the learnt models by comparing the average SSIM and PSNR of image sequences obtained by unrolling the dynamics from $Z_{0}$ using the learnt transition function $f_{\theta}$, and then generating images by passing the resulting state trajectory through the learnt decoder $r_{\phi}$. All results corresponding to the first experiment setting are averaged over $5$ independent runs, and all results corresponding to the second experiment setting are averaged over $9$ independent runs.
For the first setting, results for different configurations of the diffusion resampler are presented in Table 13. Results for different configurations of the baselines are presented in Table 14. For the second, more challenging setting, in which resampling is invoked at nearly every filtering step, the corresponding results are presented in Tables 15 and 16, respectively. Figure 8 shows the median loss evolution during training for the best model configuration (in terms of mean SSIM and PSNR) from each resampling class in each experiment setting. Figure 9 shows mean prediction SSIM/PSNR for the second setting; see Figure 5 for a comparison to the corresponding results in the first setting.
The difference in SSIM/PSNR between the two experiments indicate that the second setting is more challenging for all resampling methods. In the first setting, performance is competitive, and diffusion resampling provides a stable end-to-end optimisation and effective integration into the high-dimensional learning pipeline. Notably, resampling is not always necessary for convergence in end-to-end training using particle filters, and prior work has reported cases where omitting resampling can be more effective (see, e.g., Karkus et al., 2018). In our experiments, we consider end-to-end training with resampling, and compare resamplers while holding the rest of the system fixed. Resampling itself can interact with the optimisation process in non-trivial ways, and we thus primarily evaluate the comparison by end-to-end performance. However, tempering the observation likelihood in the first setting reduces weight concentration: the weights become more uniform and ESS remains high. With this ESS-based adaptive resampling, the resampling can therefore be triggered infrequently. As such, to further stress test differentiable resampling in this context, we consider the second setting with more concentrated weights and resampling at nearly every filtering step.
As we have seen, even in this challenging setting, diffusion resampling remains competitive among the baselines. Figure 10 shows a qualitative comparison of a top performing model in each resampling class in this setting. Here, the diffusion resampler is among the two top-performing baselines. While soft resampling occasionally shows better performance in terms of final metrics compared to diffusion resampling, see Figure 9, we note that it also exhibits significantly worse weight degeneracy compared to diffusion resampling and the other baselines. In addition, it is by construction not fully differentiable (see Table 17).
Identifiability of the latent space
@@
Because both the transition model $f_{\theta}$ and the decoder $r_{\phi}$ are learnt only through the image observation likelihood, the latent state is typically identifiable only up to (possibly nonlinear) invertible transformations, and there is no unique correspondence between the learnt states and the true physical coordinates without special constructions (see, e.g., Greydanus et al., 2019, for a similar issue). For this reason, we do not directly evaluate the pendulum dynamics in state space, but instead assess model quality in the observation space via image reconstruction metrics. A similar evaluation strategy is common in recent work on latent SDEs for high-dimensional sequence data. For example, both Bartosh et al. (2025) and Course and Nair (2023) primarily report qualitative comparisons of generated image sequences in comparable settings, rather than quantitative metrics in the latent state space, reflecting the view that high-quality reconstruction is indicative of a meaningful latent representation. Nevertheless, we find that sometimes the learnt dynamics closely match the true pendulum dynamics up to a linear transformation of the latent coordinates. Figure 11 illustrates two such examples with the latent dynamics learnt using diffusion resampling, where a simple linear transformation of the latent trajectory yields a close match to the ground-truth pendulum state (coefficient of determination $R^{2}\approx 0.98$ (left) and $R^{2}\approx 0.92$ (right), respectively). This indicates that the models have learnt meaningful latent representations of the underlying pendulum dynamics, albeit in a transformed coordinate system.
@@
Table 13: Prediction quality in terms of structural similarity index (SSIM) and PSNR (higher the better) for the pendulum dynamics tracking experiment, with respect to models trained using diffusion resampling in the first experiment setting. Each result is presented in terms of mean over 5 independent runs, together with the corresponding standard deviation, recorded after 3000 training iterations. For two configurations only 1 out of 5 (*) and 3 out of 5 (**) runs, respectively, were non-divergent.
@@
Table 14: Prediction quality (SSIM, PSNR) for the pendulum dynamics tracking experiment, with respect to models trained using OT, Gumbel and Soft resampling. Results correspond to the first experiment setting.
@@
Table 15: Prediction quality in terms of structural similarity index (SSIM) and PSNR (higher the better) for the pendulum dynamics tracking experiment, with respect to models trained using diffusion resampling in the second experiment setting. Each result is presented in terms of mean over 9 independent runs, together with the corresponding standard deviation, recorded after 1500 training iterations. For four configurations only 8 out of 9 (*) runs, respectively, were non-divergent.
Method Euler–Maruyama Jentzen–Kloeden Tweedie
SDE ODE SDE ODE SDE
SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR
$T=1$, $K=4$ $0.671^{*}\pm 0.082$ $17.2^{*}\pm 0.666$ $0.607\pm 0.091$ $17.2\pm 0.834$ $0.517^{*}\pm 0.134$ $16.9^{*}\pm 0.528$ $0.507\pm 0.048$ $16.5\pm 0.443$ $0.549^{*}\pm 0.081$ $16.7^{*}\pm 0.666$
$T=1$, $K=8$ $0.576^{*}\pm 0.080$ $16.6^{*}\pm 0.542$ $0.511\pm 0.070$ $16.7\pm 0.554$ $0.599\pm 0.101$ $17.1\pm 0.579$ $0.481\pm 0.074$ $16.5\pm 0.253$ $0.503\pm 0.067$ $16.5\pm 0.32$
$T=1$, $K=16$ $0.576\pm 0.082$ $16.6\pm 1.19$ $0.280\pm 0.112$ $16.1\pm 0.235$ $0.557\pm 0.101$ $16.9\pm 0.744$ $0.239\pm 0.082$ $16.0\pm 0.202$ $0.227\pm 0.074$ $16.1\pm 0.197$
Table 16: Prediction quality (SSIM, PSNR) for the pendulum dynamics tracking experiment, with respect to models trained using OT, Gumbel and Soft resampling. Results correspond to the second experiment setting. For OT, only 1 run (*) was non-divergent.
Method SSIM PSNR
OT ($\varepsilon=0.5$) $0.697^{*}$ $17.4^{*}$
OT ($\varepsilon=1.0$)
OT ($\varepsilon=1.5$)
Gumbel (0.1) $0.509\pm 0.060$ $16.7\pm 0.400$
Gumbel (0.3) $0.593\pm 0.083$ $16.8\pm 0.508$
Soft (0.7) $0.720\pm 0.107$ $17.7\pm 0.707$
Soft (0.9) $0.704\pm 0.114$ $17.3\pm 0.872$
Network architectures
@@
Both the neural network modelling the latent dynamics, $f_{\theta}$, and the neural network modelling the decoder, $r_{\phi}$, operate on a feature vector $\mathbf{h}(Z_{j})=\begin{bmatrix}\sin(Z_{j}^{(1)})&\cos(Z_{j}^{(1)})&Z_{j}^{(2)}\,/\,10\end{bmatrix}$, which embeds the angle and scales the velocity. While the system is initialised with physical coordinates $Z_{0}$, both networks are unconstrained and unregularised for all $t>0$. As discussed above, this means that the system is free to learn a latent manifold sufficient for image reconstruction, even if the latent representation is a non-linear transformation of the true physical coordinate system. We employ a SIREN (Sitzmann et al., 2020) architecture to model the latent dynamics. The network takes the concatenated input $\begin{bmatrix}\mathbf{h}(Z_{j})&\zeta_{j}\end{bmatrix}\in\mathbb{R}^{5}$ and outputs the state derivative. The update rule in the first setting is $Z_{j+1}=Z_{j}+\mathrm{Net}(\begin{bmatrix}\mathbf{h}(Z_{j})&\zeta_{j}\end{bmatrix})\,\Delta_{j}$. In the second setting, where we invoke resampling at nearly all filtering steps, we make an additional additive noise assumption, omit concatenating the input and let $Z_{j+1}=Z_{j}+\mathrm{Net}(\mathbf{h}(Z_{j}))\,\Delta_{j}+\zeta_{j}$. Without this assumption, the neural network is free to learn to neglect the noise, and in the first setting, we indeed found that the learnt dynamics network often had this issue. This can happen in end-to-end learning since minimising the optimisation objective does not necessarily require the network to optimise for a good filtering distribution, and reducing the randomness might, for instance, help stabilise optimisation.
In both settings, the network consists of 3 hidden layers of 256 units with sine activations ($\omega_{0}=8.0$ in the first layer). The decoder maps $\mathbf{h}(Z_{j})$ to observations and is identical in both settings. It first processes the input via two linear layers (16 and 256 units), reshaping the output into a spatial feature map of size $4\times 4\times 16$. This is followed by three transposed convolution layers (kernel size $3\times 3$, stride 2) with output channels of 32, 16, and 1, respectively, to upsample to the final $32\times 32$ image. ReLU activations are used for all hidden layers, while the output layer is linear.
Figure 8: (Left) Median loss during the training process for the top performing configuration in each resampling class in the first setting. We note that diffusion resampling achieves the lowest loss. (Right) Median loss during the training process for the top performing configuration in each resampling class in the more challenging setting. We note that for OT, the reported median loss is based on the single successful run and is therefore not statistically meaningful. While Soft achieves the lowest median loss (see the text for discussion), diffusion resampling remains competitive and yields low, stable loss, whereas Gumbel performs markedly worse than all baselines.
Figure 9: Mean prediction SSIM and PSNR for the best (by mean) model configuration of each resampler in the more challenging setting. Individual runs are shown as scatter points. We find that although Soft can occasionally give better results than diffusion resampling, its results exhibit large variability. We also observe that Gumbel performs worse than diffusion resampling.
Figure 10: Qualitative comparisons of the learnt pendulum dynamics trained end-to-end using a differentiable resampler in the SMC training loop. The results correspond to the second, more challenging experiment setting using resampling at nearly all filtering steps. The ground truth (green) is overlaid with model predictions (purple). White pixels indicate perfect alignment, while coloured regions highlight positional discrepancies (e.g., phase lag). Snapshots are shown at eight evenly spaced time points over the 4 second simulation (read from left to right and top to bottom). Each panel shows a top result in terms of mean SSIM for each differentiable resampling method in our comparison: a) Diffusion (mean SSIM/PSNR $0.761/17.0$, b) Soft ($0.816/19.2$), c) OT ($0.697/17.4$), d) Gumbel ($0.663/16.9$).
@@
Figure 11: (Left) Ground-truth pendulum state trajectories (dashed) and predictions in the original neural (solid with crosses) and transformed (solid) coordinate systems, for a top-performing model trained with diffusion resampling ($T=1$, $K=4$) in the first setting. The dynamics model is trained until convergence, and correspond to the image predictions presented in Figure 6. (Right) Same as the previous image, but for a model optimised in the more challenging setting. The corresponding image predictions are presented in Figure 10a.
Appendix K Additional related work
In addition to the discussion in Section 5, there are also connections to other, less central, lines of work. For instance, the ensemble score estimator is related to the kernel projection used in Stein variational gradient flow (Liu, 2017). The controlled gain term in Yang et al. (2013) also plays a role analogous to resampling in a continuous-time limit, but involves a computationally demanding PDE-solving step. While there are interesting connections to our work, the approach focuses on continuous-time flow filtering (cf. Kang et al., 2025) and leads to algorithms that are no longer standard SMC.
Appendix L Take-away messages
@@
  • Diffusion resampling is useful not only for differentiation, but also for reducing resampling error in general. The primary reason for this is that diffusion resampling can take additional information (e.g., $\pi_{\mathrm{ref}}$) of the target into account. When we have a complete information $\pi_{\mathrm{ref}}=\pi$, the diffusion resampling becomes an optimal resampling algorithm. Corollary 1 explicitly shows how the resampling error can be controlled depending on how well $\pi_{\mathrm{ref}}$ approximates $\pi$, calibrated by $b$. In contrast, multinomial resampling only uses information from the given samples $\{(w_{i},X_{i})\}_{i=1}^{N}$ which uses incomplete and less information compared to diffusion resampling based on first principle.
@@
  • If a good reference distribution is hard to construct, diffusion resampling may not be optimal for a pure resampling problem outside the SSM context. For instance, as shown in the Gaussian mixture experiment, the large discrepancy between the Gaussian reference and the highly non-Gaussian target may prolong the diffusion time. However, for SSMs, thanks to their sequential structure, constructing an effective reference becomes straightforward.
Table 17: Comparison among commonly used resampling schemes with their calibration parameters. The computational complexity is analysed per sample which can be embarrassingly parallelised over the samples. The complexity of OT is given by Luo et al. (2023a), where we here parametrise the Sinkhorn precision with the regularisation parameter $\varepsilon$ to unify comparison. By “fully differentiable” we mean that the pathwise gradient is well defined, for instance, the soft resampling is only partially differentiable.
@@
Diffusion $K$ Yes Yes, as $N(K)\to\infty$ No O(KN)O(K\,N)
OT $\varepsilon$ Yes Yes, as $N(\varepsilon)\to\infty$ No $O(N^{2}\log(N)^{-1}\,\varepsilon^{-1})$